If the length of DNA is 1.36 mm, the number of base pair and nucleosomes in it will be?
1
Expert's answer
2015-11-23T14:28:36-0500
Answer
36.9 L
Solution
PV = nRT
V = nRT / P
Mg(s)+2HCl(aq)→MgCl2(aq)+H2(g)
n(H2) = n (Mg) = m (Mg) / M (Mg) = 41.0 g / (24.0 g/mol) = 1.7 mol
R = 62360 mmHg/mol*K
T = 23 ˚C + 273 = 296 K
V(H2) = (1.7 mol * 62360 mmHg / mol*K * 296) / 850 mm Hg = 36917.1 ml = 36.9 L